THE MICROWAVE SPECTRUM AND MOLECULAR STRUCTURE OF TRIFLUOROETHYLENE-ACETYLENE

\underline{HELEN O. LEUNG}, MARK D. MARSHALL, WINN T. CASHION, Department of Chemistry, Amherst College, P.O. Box 5000, Amherst, MA 01002-5000.

The rotational spectra of six isotopomers of the 1,1,2–trifluoroethylene-acetylene complex have been collected in the 6–21 GHz region. In addition to the most abundant isotopomer, spectra are obtained for four isotopomers containing substitution in the HCCH subunit (HCCD, H^{13}C^{13}CH, and H^{13}CCH, which forms two different isotopomers for the complex) and for one isotopomer containing a substitution in the CHFCF_{2} subunit (CHF^{13}CF_{2}). The spectra of the three isotopomers with a single substitution of ^{13}C are observed in natural abundance upon pulsed jet expansion of a mixture of HCCH, CHFCF_{2}, and Ar. Both a- and b-type transitions are seen, and the deuterium nuclear quadrupole hyperfine structure in HCCD-CHFCF_{2} is resolved and analyzed. Both the Kraitchman substitution coordinates and the spectroscopic constants are consistent with a planar structure in which a hydrogen bond is formed between H in HCCH and the F on C-2 in CHFCF_{2} with a secondary interaction between the H (geminal to the F involved in the hydrogen bond) in CHFCF_{2} and the acetylenic bond.